OpenStudy (anonymous):

Find all soultions for sec3x=-1

OpenStudy (ddcamp):

\[\frac{ 1 }{ \cos(3x) } = -1\] Solve for cos(3x), Then use the unit circle.

OpenStudy (anonymous):

cos(3x)=-1, 3x=180, hence x=60 (of course you could add or subtract 360n)